OpenStudy (anonymous):

Front-row tickets for a concert cost $60. After every 5 rows, the ticket price goes down by $5. What is the total amount of money generated by a full house? Info: There are 20 rows that increase by 3 seats each time starting with 18.

OpenStudy (anonymous):

ive been stuck for over an hour now. Please help!

OpenStudy (anonymous):

Well, You should start by: ((60*18)*5)+((55*15)*5)+((50*12)*5)+((45*9)*5)

OpenStudy (anonymous):

Where did you get the number from? This is so confusing..

OpenStudy (anonymous):

What does it equal? I can put it in steps if you'd like to make it look simpler.

OpenStudy (anonymous):

Steps would be awesome! Thank you!

OpenStudy (anonymous):

Okay :P Step 1: ((60*18)*5)

OpenStudy (anonymous):

5400

OpenStudy (anonymous):

Step 2: 5400+((55*15)*5)

OpenStudy (anonymous):

4125.

OpenStudy (anonymous):

Then 3000

OpenStudy (anonymous):

Step 3: ((50*12)*5)

OpenStudy (anonymous):

and 2025

OpenStudy (anonymous):

Step 4: ((45*9)*5)

OpenStudy (anonymous):

How are you getting these numbers?

OpenStudy (anonymous):

From the information, the question. It's a simple process, you just have to break it down into parts.

OpenStudy (anonymous):

Could you explain in detail please? :)

OpenStudy (anonymous):

Sure. So I got the numbers from Step 1 because it told me $60 dollars for a row of tickets, which is 18 seats. The other information told me it went on for 5 rows, so in all the question by numbers would be: (60*18)*5

OpenStudy (anonymous):

I got the numbers for Step 2 from $60 (subtract $5 because it said after 5 rows to subtract $5), which would create $55. The number in each row went UP (OOPS, OMG I'M SO SORRY, I THOUGHT IT SAID DECREASE) by 3. So 18 plus 3 equals 21. Then 5 rows, so it would become: (55*21)*5

OpenStudy (anonymous):

I got the numbers from Step 3 by 55-5=50. Then, 21+3=24. Last, 5 rows of seats. (50*24)*5

OpenStudy (anonymous):

For Step 4, there's 50-5=45 dollars per ticket. Then, 24+3=27 seats per row and 5 rows of seats. (45*27)*5

OpenStudy (anonymous):

So 20 rows (5+5+5+5=20 rows) if you add those all together, you should get your answer. ((60*18)*5)+((55*21)*5)+((50*24)*5)+((45*27)*5)= YOUR ANSWER HERE

OpenStudy (anonymous):

Does it make a little bit more sense? Usually, you have to break each number down, put them in groups, and do each group at a time.

OpenStudy (anonymous):

ohhh okay! thank you.. do you know how to put something like this into an equation?

OpenStudy (anonymous):

Yeah, like a number equation? I did...this one: ((60*18)*5)+((55*21)*5)+((50*24)*5)+((45*27)*5)= YOUR ANSWER HERE Step 1: ((60*18)*5) Step 2: ((55*21)*5) Step 3: ((50*24)*5) Step 4: ((45*27)*5) Step 5: Add all of your answers from Steps 1, 2, 3, and 4. That will be your answer.

OpenStudy (anonymous):

Does that work for you? If you need any additional help, just tell me.

OpenStudy (anonymous):

thats awesome thanks for your help!